# Soliton resuscitations: asymmetric revivals of the breathing mode of an atomic bright soliton in a harmonic trap

### v3

@@ Abstract
Collective excitations of individual solitons have attracted interest as well, and indeed the distinction between collective motional modes of multi-soliton groups, and the vibration-like excitation modes of individual solitons, is not really sharp. Higher-order “breather” solitons[11, 12, 13, 14] can be considered single solitons with nonlineary large-amplitude breathing excitations, or they can be seen as interactions of two or more solitons that repeatedly merge and separate. Quantum effects in these nonlinear excitations, for the dilute condensed Bose gas beyond the mean-field nonlinear Schrödinger equation, have been studied theoretically using truncated Wigner[11], Bethe Ansatz[12] and Bogoliubov-de Gennes theory[13].
@@
Figure 2: Asymmetrical resuscitations of soliton breathing in a harmonic trap of frequency $\omega$, as shown in the peak condensate density $|\psi(0,t)|^{2}$ as a function of time in trap periods. The evolution is given by the one-dimensional Gross-Pitaevskii equation from the non-stationary initial wave function $\psi(x,0)=A(\tilde{\kappa}/\sqrt{g})\mathrm{sech}(\tilde{\kappa}x)$ for $\tilde{\kappa}=\sqrt{50}a_{0}^{-1}$, where $a_{0}=\sqrt{\hbar/(2M\omega)}$ (with atomic mass $M$) is the harmonic trap width. The case $A=1$ would be a perfect bright soliton, with no breathing; in this case $A=1.01$, providing a small-amplitude breathing excitation at the chemical potential frequency $\hbar\kappa^{2}/(2M)=A^{4}(\tilde{\kappa}a_{0})^{2}\omega\doteq 52\omega$, while the breathing amplitude revives every half-period of the harmonic trap. The first revival begins noticeably sooner than a trap half-period. Note the “trumpet-shaped” asymmetrical envelopes of the revived oscillations.
@@
To make this perturbed $\psi(x,t)$ solve the Gross-Pitaevskii equation (3) to first order in $\delta\psi$, particular functional forms are required for the mode functions $u_{k}(x)$ and $v_{k}(x)$. These forms are known explicitly and exactly for the case $V=0$ [16], and it will be a major goal in this paper to find them approximately for a harmonic $V$. For $V=0$ we have the normalized mode functions are[16]
@@
In this paper we will analyse only the particular kind of small perturbations around $\Psi_{0}$ that evolve from slightly stretched or compressed initial conditions of the form
@@
With this specific form (II.3) of $\beta_{k}$, we can integrate (VII.1) to obtain the breathing perturbation $\delta\psi(x,t)$ for all $t$. For small $\delta$, the density expanded to first order in $\delta$
@@
is an excellent approximation to the exact $|\psi(x,t)|^{2}$ over many breathing periods. A simple measure of the overall amplitude of the breathing mode is then $|\psi(0,t)|^{2}$, which from (VII.1), (II.2), and (II.3) is given by
@@
Because the Bogoliubov-de Gennes breathing oscillation dies away from dispersion, at very late times the post-Bogoliubov-de Gennes $\mathcal{O}(\delta^{2})$ corrections to $\psi$ may be more significant locally than the leading-order term (13). We can actually identify the main post-Bogoliubov-de Gennes effect, however, using only the leading order $\delta\psi(x,t)$ from (VII.1) with the coefficients $\beta_{k}$ from (II.3).
@@ III.3.1 Discrete frequencies
Equation (25 says that $\Delta_{n}$ rises smoothly from 0 for small $n$ through 1 around $n\simeq\varepsilon^{-2}$, and ultimately approaches 2 for $n\to\infty$. The fact that $\Delta_{n\to\infty}=2$ means that $\nu_{N-2}=N$ for $N\to\infty$, so that the same large frequency range which holds $N\gg 1$ harmonic oscillator eigenfrequencies only holds N2N-2 of the bright soliton Bogoliubov-de Gennes eigenfrequencies.
@@
This subtle distortion of the Bogolibov-de Gennes eigenspectrum, in comparison to the evenly spaced harmonic oscillator spectrum, will be responsible for the main effect that we discuss in this paper. The fact that all the even-mode eigenfrequencies $\Omega_{2n}$ are spaced nearly $2\omega$ apart is responsible for the revivals of soliton breathing after every half-period $\pi/\omega$. The fact that all the intervals between successive $\Omega_{2n}$ are actually not quite exactly $2\omega$, but instead are slightly and systematically stretched according to (25) and (24), is responsible for the increasingly assymetrical forms of the resuscitations that we saw in Fig. 2.
@@ III.3.2 Normalization
to approximate (IV.1) as the $B_{0}(t)$ (II.3) that we found for the breathing amplitude of the untrapped soliton:
@@
Do the terms in the sum (IV.1) actually vary slowly with $n$, though? Whenever $n$ appears in the summand (IV.1), it is indeed multiplied by the small factor $\varepsilon^{2}$—except for the one term $2i\omega tn$ that appears in the exponent. For times much less than the trap period, this term also varies slowly with $n$, and so for $t\ll 1/\omega$ the trapped bright soliton still breathes just as it does without any trap, and the breathing amplitude decays in a $1/t$ envelope. For times $t\gtrsim 1/\omega$, however, the sum (IV.1) is not well approximated by the integral (34). The non-Markovian nature of the Trap Zone as an environment for the breathing soliton becomes apparent.
For general times of order $t\sim 1/\omega$, the sum (IV.1) has no simple evaluation. We can consider times close to a whole number of half-periods, however,
@@
for whole-number $N$ and $|\Delta t|\ll 1/\omega$. For $N\gtrsim\varepsilon^{-2}$, $2N\cos^{-1}\big{(}(1-2\varepsilon^{2}n)/(1+2\varepsilon^{2}n)\big{)}$ does not vary slowly enough with $n$ for any integral approximation to be valid, but as long as $N\ll\varepsilon^{-2}$, the summand in (IV.1) again varies slowly with $n$, because $e^{2i\pi Nn}=1$. This yields the modified integral
@@
$B_{N}(\Delta t)$ $=2\mathrm{Re}\left[e^{i\frac{N\pi}{2}}e^{i\frac{\hbar\kappa^{2}}{2M}\Delta t}\int_{0}^{\infty}\!dz\,\mathrm{sech}\Big{(}\frac{\pi z}{2}\Big{)}e^{i\tau z^{2}}\left(\frac{1+iz}{1-iz}\right)^{2N}\right]\;,$ (37)
@@
After Eqns. (23-25) for the collective mode spectrum, Eqn. (37) for the resuscitation profiles in time is our paper’s final major result.
For $N=0$, $B_{0}(\Delta t)=B_{0}(-\Delta t)$, since $\Delta t\to-\Delta t$ just changes the term in square brackets in (37) to its complex conjugate and this does not affect its real part. For $N>0$, however, this no longer true, and $B_{N>0}(\Delta t)$ is asymmetric in $\Delta t$, with the asymmetry becoming more marked with increasing $N$. Figure 4 shows a direct comparison between $g|\psi(0,t)|^{2}/\tilde{\kappa}^{2}$ from Fig. 2 and the Bogoliubov-de Gennes approximations $(1+\delta)^{2}[1-\delta B_{N}(t-N\pi/\omega)]$ in intervals around three successive resuscitation times $t=N\pi/\omega$. Since $\varepsilon\doteq 0.14$ in this case is not actually very small, the integral approximation becomes poorer at larger $\Delta t$, and for later resuscitations it evidently exaggerates the severity of the post-revival drop in breathing amplitude. The agreement between nonlinear mean field theory and the Bogoliubov-de Gennes approximation is nevertheless close enough to confirm that the asymmetric profile of the resuscitations is indeed due to the slight non-uniformity of the collective mode frequency spectrum.

### v4

@@ Abstract
Collective excitations of individual solitons have attracted interest as well, and indeed the distinction between collective motional modes of multi-soliton groups, and the vibration-like excitation modes of individual solitons, is not always sharp, because in some cases the multiple solitons never fully separate from each other. Higher-order “breather” solitons[11, 12, 13, 14] can be considered single solitons with nonlinearly large-amplitude breathing excitations, or they can be seen as interactions of two or more solitons that repeatedly merge and (perhaps only partially) separate. Quantum effects in these nonlinear excitations, for the dilute condensed Bose gas beyond the mean-field nonlinear Schrödinger equation, have been studied theoretically using truncated Wigner[11], Bethe Ansatz[12] and Bogoliubov-de Gennes theory[13].
@@
Figure 2: Asymmetrical resuscitations of soliton breathing in a harmonic trap of frequency $\omega$, as shown in the peak condensate density $|\psi(0,t)|^{2}$ as a function of time in trap periods. The evolution is given by the one-dimensional Gross-Pitaevskii equation from the non-stationary initial wave function $\psi(x,0)=A(\tilde{\kappa}/\sqrt{g})\mathrm{sech}(\tilde{\kappa}x)$ for $\tilde{\kappa}=\sqrt{50}a_{0}^{-1}$, where $a_{0}=\sqrt{\hbar/(2M\omega)}$ (with atomic mass $M$) is the harmonic trap width. The case $A=1$ would be a perfect bright soliton, with no breathing; in this case $A=1.01$, providing a small-amplitude breathing excitation at the chemical potential frequency $\hbar\kappa^{2}/(2M)=A^{4}(\tilde{\kappa}a_{0})^{2}\omega\doteq 52\omega$. The amplitude initially decays in an envelope that closely fits the $t^{-1/4}$ steepest descents approximation of Eqn. (15) (dashed curves), but then the breathing amplitude revives: it recovers to a maximum amplitude approximately every half-period of the harmonic trap. In fact the first revival begins noticeably sooner than a trap half-period; note as well the “trumpet-shaped” asymmetrical envelopes of the revived oscillations.
@@
To make this perturbed $\psi(x,t)$ solve the Gross-Pitaevskii equation (3) to first order in $\delta\psi$, particular functional forms are required for the mode functions $u_{k}(x)$ and $v_{k}(x)$. These forms are known explicitly and exactly for the case $V=0$ [16], and it will be a major goal in this paper to find them approximately for a harmonic $V$. For $V=0$, we have the normalized mode functions that are[16]
@@
The spectrum of $k$ is indeed continuous, and no $u_{k}$ actually decay at large $|x|$: the bright soliton does not have a discrete spectrum of collective multipole excitation modes, as a condensate held in an external trap does. Long-wavelength perturbations, with amplitude concentrated in $|k|\ll\kappa$, will translate the soliton, if they have odd spatial parity; if they have even parity, they can make the soliton “breathe”. Otherwise, if there are initial perturbations present that change the condensate profile away from the perfect soliton’s $\mathrm{sech}$ profile in more complicated ways than translation or breathing, then these excitations will have to have significant amplitude at $|k|\gtrsim\kappa$, in order to affect $\psi(x,0)$ on wavelengths shorter than the soliton width. All such initial excitations will immediately disperse away from the soliton at speeds of order $\hbar\kappa/M$: they represent pure transients, and not longer-lived collective excitations of the soliton. So, apart from translation, breathing is actually the only thing like a collective mode which a bright soliton can possess.
In this paper, therefore, we will analyse only the particular kind of small perturbations around $\Psi_{0}$ that evolve from slightly stretched or compressed initial conditions of the form
@@
With this specific form (II.3) of $\beta_{k}$, we can integrate (II.2) to obtain the breathing perturbation $\delta\psi(x,t)$ for all $t$. For small $\delta$, the density expanded to first order in $\delta$
@@
is an excellent approximation to the exact $|\psi(x,t)|^{2}$ over many breathing periods. A simple measure of the overall amplitude of the breathing mode is then $|\psi(0,t)|^{2}$, which from (II.2), (II.2), and (II.3) is given by
@@
Because the Bogoliubov-de Gennes breathing oscillation dies away from dispersion, at very late times the post-Bogoliubov-de Gennes $\mathcal{O}(\delta^{2})$ corrections to $\psi$ may be more significant locally than the leading-order term (13). We can actually identify the main post-Bogoliubov-de Gennes effect, however, using only the leading order $\delta\psi(x,t)$ from (II.2) with the coefficients $\beta_{k}$ from (II.3).
@@ III.3.1 Discrete frequencies
Equation (25) says that $\Delta_{n}$ rises smoothly from 0 for small $n$ through 1 around $n\simeq\varepsilon^{-2}$, and ultimately approaches 2 for $n\to\infty$. The fact that $\Delta_{n\to\infty}=2$ means that $\nu_{N-2}=N$ for $N\to\infty$, so that the same large frequency range which holds $N\gg 1$ harmonic oscillator eigenfrequencies only holds N2N-2 of the bright soliton Bogoliubov-de Gennes eigenfrequencies.
@@
This subtle distortion of the Bogolibov-de Gennes eigenspectrum, in comparison to the evenly spaced harmonic oscillator spectrum, will be responsible for the main effect that we discuss in this paper. The fact that all the even-mode eigenfrequencies $\Omega_{2n}$ are spaced nearly $2\omega$ apart is responsible for the revivals of soliton breathing after every half-period $\pi/\omega$. The fact that all the intervals between successive $\Omega_{2n}$ are actually not quite exactly $2\omega$, but instead are slightly and systematically stretched according to (25) and (24), is responsible for the increasingly asymmetrical forms of the resuscitations that we saw in Fig. 2.
@@ IV.1.1 Evaluating $B(t)$ for small $t\ll\omega^{-1}$
to approximate (IV.1.1) as the $B_{0}(t)$ (II.3) that we found for the breathing amplitude of the untrapped soliton:
@@
Do the terms in the sum (IV.1.1) actually vary slowly with $n$, though? Whenever $n$ appears in the summand (IV.1.1), it is indeed multiplied by the small factor $\varepsilon^{2}$—except for the one term $2i\omega tn$ that appears in the exponent. For times much less than the trap period, this term also varies slowly with $n$, and so for $t\ll 1/\omega$ the trapped bright soliton still breathes just as it does without any trap, and the breathing amplitude decays in a $1/t$ envelope. For times $t\gtrsim 1/\omega$, however, the sum (IV.1.1) is not well approximated by the integral (34). The non-Markovian nature of the Trap Zone as an environment for the breathing soliton becomes apparent.
IV.2 IV.B Evaluating $B(t)$ for $t=N\pi/\omega+\Delta t$
For general times of order $t\sim 1/\omega$, the sum (IV.1.1) has no simple evaluation. We can consider times close to a whole number of half-periods, however,
@@
for whole-number $N$ and $|\Delta t|\ll 1/\omega$. For $N\gtrsim\varepsilon^{-2}$, $2N\cos^{-1}\big((1-2\varepsilon^{2}n)/(1+2\varepsilon^{2}n)\big)$ does not vary slowly enough with $n$ for any integral approximation to be valid, but as long as $N\ll\varepsilon^{-2}$, the summand in (IV.1.1) again varies slowly with $n$, because $e^{2i\pi Nn}=1$. This yields the modified integral
@@
$B_{N}(\Delta t)$ $=2\mathrm{Re}\left[e^{iN\pi\varepsilon^{-2}}e^{i\frac{N\pi}{2}}e^{i\frac{\hbar\kappa^{2}}{2M}\Delta t}\int_{0}^{\infty}\!dz\,\mathrm{sech}\Big(\frac{\pi z}{2}\Big)e^{i\frac{\hbar\kappa^{2}}{2M}\Delta t\,z^{2}}\left(\frac{1+iz}{1-iz}\right)^{2N}\right]\;,$ (37)
@@
After Eqns. (23-25) for the collective mode spectrum, Eqn. (37) for the resuscitation profiles in time is our paper’s final major result. What does it tell us?
IV.3 IV.C Features of the resuscitation episodes $B_{N}(\Delta t)$
Equation (37) defines the oscillation amplitude of the soliton’s peak density for a range of time $\Delta t$ around the $N$th revival time, $t=N\pi/\omega+\Delta t$ as defined in (35). Even before we attempt to evaluate the integral over $z$ in (37), we can see that the two relevant frequency scales $\omega$ and $\hbar\kappa^{2}/(2M)$ appear in $B_{N}(\Delta t)$ in quite separate roles.
IV.3.1 The phase shift $N\pi/\varepsilon^{2}$
First of all, the trap frequency $\omega$ only appears in (37) in one single place: it is inside the large factor
$\varepsilon^{-2}=\frac{\hbar\kappa^{2}}{2M\omega}$ (39)
which appears in the first phase in $B_{N}(\Delta t)$. In our small-$\varepsilon$ limit this phase is enormous for $N\geq 1$.
This large phase must indeed really be present, since the number of breathing mode periods in one-half of a trap period is a large number that does not have to be an integer. It is not easy, however, to make a reliable prediction of the value of this huge phase modulo $2\pi$. All of our calculations have neglected higher orders of $\varepsilon$, and if any of these neglected corrections turn out to shift the overall phase of the breathing mode, their effects on the phase may well be as great as $N\pi\varepsilon^{-2}$ modulo $2\pi$, even if they are much smaller than $N\pi\varepsilon^{-2}$.
Even if we had been able to perform our Bogoliubov-de Gennes calculations to all orders in $\varepsilon$, furthermore, it would still be beyond the scope of this paper to assess possible additional secular phase drifts due to nonlinear mean-field effects, or quantum correction to mean-field physics, or limitations of the quasi-one-dimensional approximation, all of which could likewise shift the breathing mode phase over a large number of breathing periods. And, on the other hand, even if we could compute all these effects, it is unlikely that the soliton chemical potential $\hbar^{2}\kappa^{2}/(2M)$ can experimentally be controlled precisely enough to make $N\pi\varepsilon^{-2}$ modulo $2\pi$ a reproducible phase instead of an effectively random one.
The phase factor $e^{iN\pi\varepsilon^{-2}}$ in (37) must probably therefore be considered uncertain, or even randomly varying in each experimental run. All the other features of $B_{N}(\Delta t)$—the frequency of the breathing oscillations, and their asymmetrically reviving and decaying amplitude envelope—should in contrast be robust predictions of mean field theory, as linearized for a small-amplitude collective excitation.
@@ IV.3.1 The phase shift $N\pi/\varepsilon^{2}$
IV.3.2 Independence of $\omega$
Unlike the doubtful $e^{iN\pi\varepsilon^{-2}}$ phase factor, the other features of $B_{N}(\Delta t)$ are all completely independent of the trap frequency $\omega$, as long as we are indeed in the $\omega\ll\hbar\kappa^{2}/M$ limit that we have always assumed in this paper. In all cases in this regime, the soliton’s breathing oscillation always has the same frequency scale $\hbar\kappa^{2}/(2M)$ as the untrapped soliton. As we will see, the amplitude envelope around each resuscitation episode has features on the additional time scales $\hbar\kappa^{2}/(MN)$ and $\hbar\kappa^{2}/(MN^{2})$—but neither of these additional time scales involves the trap frequency $\omega$.
If we compare soliton breathing in traps of different $\omega$, therefore, we will have to wait different lengths of time between resuscitations, because the time between resuscitations is half the trap period, $\pi/\omega$. Once the resuscitation episodes begin in each trap, however, they will look the same in both cases, in spite of the different $\omega$. In this sense $B_{N}(\Delta t)$ should be considered independent of $\omega$, and universal for all $(2M\omega)/(\hbar\kappa^{2})=\varepsilon^{2}\ll 1$ (once we interpret the huge phase $e^{iN\pi\varepsilon^{-2}}$ as effectively random modulo $2\pi$).
IV.3.3 Form of $B_{N}(\Delta t)$ for smaller NN
For $N=0$, $B_{0}(\Delta t)=B_{0}(-\Delta t)$, since $\Delta t\to-\Delta t$ just changes the term in square brackets in (37) to its complex conjugate and this does not affect its real part. For $N>0$, however, this no longer true, and $B_{N>0}(\Delta t)$ is asymmetric in $\Delta t$, with the asymmetry becoming more marked with increasing $N$. Figure 4 shows a direct comparison between $g|\psi(0,t)|^{2}/\tilde{\kappa}^{2}$ from Fig. 2 and the Bogoliubov-de Gennes approximations $(1+\delta)^{2}[1-\delta B_{N}(t-N\pi/\omega)]$ in intervals around three successive resuscitation times $t=N\pi/\omega$. Since $\varepsilon\doteq 0.14$ in this case is not actually very small, the integral approximation becomes poorer at larger $\Delta t$, and for later resuscitations it evidently exaggerates the severity of the post-revival drop in breathing amplitude.
The agreement between nonlinear mean field theory and the Bogoliubov-de Gennes approximation is nevertheless close enough to confirm that the asymmetric profile of the resuscitations is indeed due to the slight non-uniformity of the collective mode frequency spectrum. Somewhat surprisingly, even the suspect phase $e^{iN\pi\varepsilon^{-2}}$ in $B_{N}(\Delta t)$ matches the nonlinear evolution very well; perhaps it helps that in this case $\varepsilon^{-2}\doteq 52$ happens to be very close to an even number.
The plots of $B_{N}(\Delta t)$ for $N\leq 3$ shown in Fig. 4 represent numerical evaluation of the $z$ integral in (37). To see what happens for larger $N$, we can pursue analytical approximations to (37) instead.
IV.3.4 Evaluating $B_{N}(\Delta t)$ for $1\ll N\ll\varepsilon^{-2}$
While our figures have shown no more than three revivals, we can see that the asymmetry persists over many revivals, by applying the saddlepoint approximation to the integral over $z$ in (37), for large $\hbar\kappa^{2}t/M$ and large $N$ (as long as $N\ll\varepsilon^{-2}$ so that the integral approximation (37) to the sum (IV.1.1) remains good). The saddlepoint condition is the cubic equation
$z_{s}^{3}+z_{s}+\frac{4NM}{\hbar\kappa^{2}\Delta t}=0\;.$ (40)
For $\Delta t>0$ there are no saddlepoints $z_{s}$ near the positive real axis of $z$ which is our integration range, and the integral is dominated by its $z=0$ endpoint. For small $z$ and large $N$, the factor $((1+iz)/(1-iz))^{2N}\doteq e^{4iNz}$, and the $z$ integral in (37) becomes close to $i/(4N)$, leaving
$B_{N}(\Delta)\xrightarrow[\Delta t>0]{}\frac{1}{2N}\cos\Big(\frac{N+1}{2}\pi+\frac{N\pi}{\varepsilon^{2}}+\frac{\hbar\kappa^{2}}{2M}t\Big)\;.$ (41)
Thus, immediately after every $N$th trap half-period, the breathing amplitude drops to $\delta/(2N)$—and then remains constant for $\mathcal{O}(16N^{2})$ breathing periods, before eventually beginning to attenuate as $|\Delta t|^{-1/2}$ (if $\varepsilon^{2}$ is small enough that the next resuscitation has not yet begun by this point). This simple behavior for $\Delta t>0$ appears at leading order in $1/(4N)$, and so we can already see it beginning in the right-most plot in Fig. 4, at just $N=3$.
For $\Delta t<0$, in contrast, the $z$ integrand in (37) has a saddlepoint $z_{s}$ on the positive real axis of $z$. The location of the saddlepoint depends on both $\Delta t$ and $N$, so that the oscillation amplitude $B_{N}(\Delta t)$ has a non-trivial envelope for negative $\Delta t$, unlike the constant amplitude $\propto 1/N$ for positive $\Delta t$. The saddlepoint approximation is remarkably close to the numerical $B_{N}(\Delta t)$ as soon as $\hbar\kappa^{2}\Delta t/M\gg 1$, but it is a complicated expression unless $\hbar\kappa^{2}\Delta t/M\gg N$, so that the cubic solution for $z_{s}$ reaches the simple limit
$z_{s}\longrightarrow-\frac{4NM}{\hbar\kappa^{2}\Delta t}\ll 1\;.$ (42)
In this small-$z_{s}$ but large-$\hbar\kappa^{2}t/M$ limit, the $z$ integral in (37) simplifies and we obtain
$B_{N}(\Delta t)\xrightarrow[\Delta t<0]{}\sqrt{\frac{\pi M}{-\hbar^{2}\kappa^{2}\Delta t}}\mathrm{Re}\left(e^{iN\pi\varepsilon^{-2}}e^{iN\frac{\pi}{2}}e^{i\frac{\hbar\kappa^{2}\Delta t}{2M}}e^{-8iN^{2}\frac{M}{\hbar\kappa^{2}\Delta t}}\Big[1+\mathrm{erf}\Big((1+i)\sqrt{\frac{4N^{2}M}{-\hbar\kappa^{2}\Delta t}}\Big)\Big]\right)\;,$ (43)
where $\mathrm{erf}(z)$ is the error function. In stark contrast to the result (41) for $\Delta t>0$, for $\Delta t<0$ the breathing amplitude $B_{N}(\Delta t)$ has a rather complicated modulating envelope that can extend over very many breathing periods. The dramatic asymmetry of the resuscitations of trapped soliton breathing modes can thus indeed persist over many revivals.